OpenStudy (anonymous):

Find the limit and confirm using the Sandwich Theorem: lim x->infinity (1 - cosx) / (x^2)

OpenStudy (anonymous):

Looked at my graphing calculator, then ate a sandwich. j/k

OpenStudy (anonymous):

I'd start with L'Hopital's rule to find the limit.

OpenStudy (s3a):

With these sandwich theorem (aka squeeze theorem) problems, you just need to take the trig function and work from there as follows.: -1 <= cos(x) <= 1 -1/x^2 <= cos(x)/x^2 <= 1/x^2 Plugging in infinity into x for the leftmost and rightmost parts of the inequality: 0 <= cos(x)/x^2 <= 0 and we can therefore conclude that the limit as x approaches infinity of cos(x)/x^2 is 0.